OpenStudy (anonymous):

is the anyone know how can i rewrite this f(x)=x^2−2x+3 in the form of f(x)=a(x-h)^2+k

OpenStudy (popitree):

\[1(x-1)^{2} + 1\]

jhonyy9 (jhonyy9):

@popitree i think this not will add to you than calcule the needed f(x) = x^2 -2x +3 try this f(x)= 1(x-1)^2 +2